OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through the points (-4,2) and (1,-1).

OpenStudy (mathstudent55):

Slope intercept form is \(y = mx + b\)

OpenStudy (mathstudent55):

m = slope b = y-intercept

OpenStudy (mathstudent55):

Since you are given two points on the line, you can find the slope.

OpenStudy (mathstudent55):

The slope, \(m\), of the line that passes through points \((x_1, ~y_1)\) and \((x_2, ~y_2) \) is \(m = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (anonymous):

I tried to solve it and I got Y=-12x-10. But that's not the right answer.

OpenStudy (mathstudent55):

What did you get for the slope?

OpenStudy (anonymous):

m=-3

OpenStudy (mathstudent55):

Let's do the slope: \(m = \dfrac{y_2 - y_1}{x_2 - x_1}\) \(m = \dfrac{-1 - 2}{1 - (-4) } = \dfrac{-3}{5} = -\dfrac{3}{5}\) This is what I get for slope.

OpenStudy (anonymous):

Your right, I messed up. I put 1-4 instead of 1-(-4). Would the answer be y=-\frac{ 12 }{ 5 }x-\frac{ 2 }{ 5 }

OpenStudy (anonymous):

I mean y=12/5x-2/5

OpenStudy (mathstudent55):

Ok, now that we know the slope is -3/5, we can replace m with -3/5: \(y = -\dfrac{3}{5}x + b\) Now all we need to do is find b. To do that just substitute x and y with the x- and y-coordinates of one known point and solve for b. We know two points, (-4, 2) and (1, -1). Let's use the fist point, (-4, 2). \(y = -\dfrac{3}{5} x + b\) \(2 = -\dfrac{3}{5} \times (-4) + b\) \(2 = \dfrac{12}{5} + b\) \(2 - \dfrac{12}{5} = b\) \(\dfrac{10}{5} - \dfrac{12}{5} = b\) \(- \dfrac{2}{5} = b\) Now that we know b, we can write the slope-intercept form using the values we have for m and b: \(y = -\dfrac{3}{5}x - \dfrac{2}{5} \)